Complete the table for the two companies using the information given below.

Company P

  Earnings

€ 6000000

Retained earnings

€ 3500000

Distributed earnings

€ 2500000

Number of shares

10000000

Current price of share

€ 5.00

Dividend per share ?

Yield ?

Earnings per share (EPS) ?

Price/ Earnings (P/E) ratio ?

 

Company Q

  Earnings

€ 6000000

Retained earnings

€ 5000000

Distributed earnings

€ 1000000

Number of shares

5000000

Current price of share

€ 6.00

Dividend per share

Yield ?

Earnings per share (EPS) ?

Price/ Earnings (P/E) ratio ?


1
Expert's answer
2021-05-07T17:26:39-0400

Let's present the calculations in EXCEL and on formulas:




"Dividend per share=\frac{Distributed earnings}{Number of shares}"


"Yield =\frac{Dividend }{Current price of share}\times100"


"Earnings per share=\frac{Retained earnings}{Number of shares}"


"Price/ Earnings (P/E) ratio=\frac{Current price of share\times Number of shares}{Retained earnings}"


Company P:

"Dividend per share=\frac{Distributed earnings}{Number of shares}=\frac{2 500 000}{ 10 000 000}=0.25"


"Yield =\frac{Dividend }{Current price of share}\times100=\frac{0.25}{5}\times100=5"


"Earnings per share=\frac{Retained earnings}{Number of shares}\frac{3 500 000}{10 000 000}=0.35"


"Price/ Earnings (P/E) ratio=\frac{Current price of share\times Number of shares}{Retained earnings}=\frac{5\times 10 000 000}{3 500 000}=14.29"


similar for the company Q